OpenStudy (anonymous):

Which expression is NOT equivalent to: 2xyz/8xz A) y/4 B) 2xyz+1/8xz+1 C) 2y/8 D) 4xy^2z/16xyz

OpenStudy (kittiwitti1):

\[\frac{2xyz}{8xz}\]What variables can cross-simplify? x and z.\[\text{So: }\frac{2y}{8}=\frac{y}{4}\]So A is a no. \[\text{Now simplify the other equations and see if they end up with }\frac{y}{4}\]

OpenStudy (anonymous):

@steve816

OpenStudy (anonymous):

Would it be B???

OpenStudy (kittiwitti1):

Did you try simplifying all the other equations?